# Rethinking of Feature Interaction for Multi-task Learning on Dense Prediction

### v1

@@ Abstract
Existing works generally adopt the encoder-decoder structure for Multi-task Dense Prediction, where the encoder extracts the task-generic features, and multiple decoders generate task-specific features for predictions. We observe that low-level representations with rich details and high-level representations with abundant task information are not both involved in the multi-task interaction process. Additionally, low-quality and low-efficiency issues also exist in current multi-task learning architectures. In this work, we propose to learn a comprehensive intermediate feature globally from both task-generic and task-specific features, we reveal an important fact that this intermediate feature, namely the bridge feature, is a good solution to the above issues. Based on this, we propose a novel Bridge-Feature-Centirc Interaction (BRFI) method. A Bridge Feature Extractor (BFE) is designed for the generation of strong bridge features and Task Pattern Propagation (TPP) is applied to ensure high-quality task interaction participants. Then a Task-Feature Refiner (TFR) is developed to refine final task predictions with the well-learned knowledge from the bridge features. Extensive experiments are conducted on NYUD-v2 and PASCAL Context benchmarks, and the superior performance shows the proposed architecture is effective and powerful in promoting different dense prediction tasks simultaneously.
@@
Multi-task Learning, Dense Prediction, Attention in Vision.
@@ 1 Introduction
Figure 1: Illustration of multi-task interaction strategies. (a) Encoder-focused [29, 23, 14, 27, 36, 15, 2]: the task-specific features directly interact with the task-generic features from the common backbone. The task-generic features contain rich low-level representation but lack of high-level representations. The interaction complexity is $O\left(n\right)$. (b) Decoder-focused [48, 9, 3, 53, 55, 50]: the task-specific features are firstly produced by deep supervision, and then interactions are conducted only based on task-specific features, which has discriminative high-level representations, but low-level representations are absent. The pair-wise interaction complexity is $O\left(n^{2}\right)$. (c) BFCI (ours): the bridge features are produced from both task-generic and task-specific features and have both rich high-level and low-level representations. The subsequent interactions based on bridge features only have $O\left(n\right)$ complexity. On the right part, we show some examples of representations from different levels in the feature map: the low-level representations contain less discriminative task information but are rich in details like object boundaries and partitions, on the contrary, the high-level representations are smooth with less details, but highly correlated to the task properties, like the highlighted floor with semantic, or areas with larger distance from the depth camera.
@@
According to where task interactions happen, existing multi-task works are roughly divided into encoder-focused and decoder-focused methods [37]. In particular, the encoder-focused methods conduct interactions of different tasks in the encoding stage, as shown in Fig. 1 (a), task-specific features interact directly with task-generic features produced by the shared backbone, to gain knowledge from multi-task shared representations and thus achieve indirect cross-task interactions. Since the image backbone/encoder is usually initialized with pre-trained parameters that contain abundant prior knowledge, and is supervised by gradients from all tasks during the training, it can produce stronger representation with rich low-level details (e.g. boundary and grouping information). For pixel-level dense prediction tasks like semantic segmentation, these representations are essential in gaining pixel relations and forming accurate semantic masks. However, in multi-task learning situations, these encoder features are task-generic, which means they only contain general dense prediction patterns, but lack distinguishable high-level representations related to each specific task. Directly conducting interactions based on task-generic features is sub-optimal, since the mining of discriminative task-specific features is insufficient and consequently limits the role of cross-task interaction [37].
@@
To alleviate the above issue, the decoder-focused methods conducting interactions in the decoding stage have been developed, where discriminative information from ground-truth labels is perceived to facilitate interactions. Specifically, the initial task predictions are first produced by preliminary decoders with deep supervision in the early decoding stage [48, 53, 55, 38, 50], aiming to decouple the generic encoder features and discover task-specific representations according to the corresponding labels. Subsequently, the cross-task interactions are conducted based on task-specific features by modules such as multi-modal distillation [48] or the ATRC [3]. These methods try to model task-pair relations, as shown in Fig. 1 (b), by extracting and transferring useful task knowledge from one task to another based on task-specific features, most of them achieve better performance compared with encoder-focused methods.
@@
i) The cross-task interactions are suffering from incompleteness and low-quality participants. Firstly, The interaction lacks essential low-level representations with rich multi-task details. In prior works, either task-generic features or task-specific features are exploited for interaction. Though beneficial task-specific features with high-level representations are gained in the initial decoding stage, the essential low-level representations gradually disappear since they are not directly required by supervision signals. However, in the following task interaction stage, only task-specific features are involved, and low-level representations that serve as an important multi-task prior are absent, leading to incomplete task interactions and limiting the model performance. Secondly, we observe that the task-specific features produced by the preliminary decoders are usually of low quality, which negatively affects the subsequent interactions, as shown in Fig 4. Since the task patterns and representations in the task-generic features are usually highly entangled, and different tasks have significantly different high-level optimizing targets, simply imposing deep supervision on preliminary decoders is not able to solve the task-pattern entanglement issue. Thus a more effective way of decoupling task-specific features from the task-generic features is required to ensure the gaining of discriminative and disentangled task representations.
@@
ii) The interaction way is inefficient with high extension costs. Here exemplified by Fig. 1 (b). Since previous decoder-focused models need to consider pair-wise task relations [3, 50], with the task number increasing, the complexity of task interactions will increase in $O\left(n^{2}\right)$, which limits the extension of previous methods on different task settings.
@@
Figure 2: The overview of our proposed method. We take depth estimation and semantic segmentation as examples. Following the basic structure of decoder-focused models, we use a shared image backbone to encode task-generic features, and a set of preliminary decoders with deep supervision are used for task-specific feature extraction. Differently, a TPP is added before the initial predictions are formed to tackle the entanglement issue. The produced task-specific features are not used for interaction directly, they are firstly processed by BFE at each scale to gain low-level representations and form the bridge features. Then, the multi-scale bridge features are used to refine task-specific features by TFR. The outputs of TFR are aggregated and fed into task-specific heads for final predictions. The detailed structures of BFE (Sec. 3.2, Fig. 3), TPP (Sec. 3.3, Fig. 5 (a)) and TFR (Sec. 3.4, Fig. 5 (b)) will be given later.
@@
To this end, we focus on improving the interaction quality for superior Multi-task Learning from the above three aspects, and propose a novel framework, namely the Bridge-Feature-Centric Interaction (BFCI) model. As shown in Fig. 1 (c), to solve the first challenge, we propose to extract and take advantage of a new type of intermediate feature that contains both rich low-level and high-level representations, namely bridge features, to ensure the integrity of feature interactions. We also propose to disentangle task patterns by jointly learning and assigning task patterns to corresponding tasks, in order to produce high-quality task-specific features for interactions. To solve the second challenge, we propose to conduct interactions directly between the bridge features and each task-specific feature, which only involves $O\left(n\right)$ complexity of task interactions. Specifically, our proposed method consists of a shared encoder backbone for the task-generic features production, the preliminary decoders with Task Pattern Propagation (TPP) in the early decoding stage producing high-quality task-specific features and handling the task-pattern-entanglement issue. The specially designed Bridge Feature Extractor (BFE), which has a transformer-based structure with cross-attention, globally queries task-specific features with the task-generic features, in order to select out high-level representations and inject them into the task-generic features with rich low-level representations, to produce multi-scale bridge features. The extracted powerful bridge features contain both high-level task-specific and low-level task-generic representations, and fertilize the formation of final predictions by the Task Feature Refiner (TFR). The overall architecture is called Bridge-Feature-Centric Interaction (BFCI).
@@
    We revisit the interacting strategies of existing multi-task dense prediction architectures, and reveal the importance of bridge features, which contain both rich low-level and high-level representations. The designed architecture, namely BFCI, performs a comprehensive interaction based on strong and representative bridge features. To the best of our knowledge, bridge features are extracted and involved in the multi-task dense prediction for the first time, which is utilized to transfer cross-task knowledge and refine the final predictions.
@@
    A transformer-based Bridge Feature Extractor (BFE) is designed to extract bridge features from both task-generic and task-specific features, and a Task Pattern Propagation (TPP) module is applied to avoid the task-pattern-entanglement issue and obtain high-quality task-specific representations. Subsequently, a Task Feature Refiner (TFR) is applied to take advantage of bridge features and refine the final task predictions.
@@ 2 Related Works
Multi-Task Learning Architectures: With the rapid development of deep CNNs, a lot of MTL works have achieved promising results [29, 14, 23, 48, 52, 53, 55, 27, 36, 38, 2, 33, 15, 20, 51, 46, 17, 50]. As mentioned in Sec. 1, existing MTL works are roughly divided into encoder-focused and decoder-focused architectures. In addition, the sharing ways of network parameters are categorized into hard and soft parameter sharing [37, 9], that correspond to directly shared parameters or indirectly constrained by regularization or losses, respectively. The encoder-focused models in [14, 33, 29] have specifically designed shared structures to explore backbone task-generic features in the encoding stage, while the decoder-focused models in [37] adopt hard parameters sharing in the encoding stage to reduce redundant parameters and FLOPs, and specially focus on the interaction among different task-specific features in decoding stages. In this paper, our proposed method is based on the decoder-focused architecture, however, also take advantage of the encoder-focused models, i.e. the exploring of task-generic features with rich low-level representations, and computational-friendly multi-task interaction costs.
@@
Task Interaction Strategies in MTL: In MTL, the learned representation from one task may be beneficial for other tasks, so it is essential to design task interaction manners to promote mutual performances, and avoid potential information inconsistencies. In the field of knowledge distillation, a multi-task student network is distilled through several single-task teacher networks, which transfer multi-task knowledge to the student network and achieve good performance [51, 20]. However, this network requires to pretrain a group of teacher networks which costs a lot by increasing the number of tasks. The typical encoder-focused models [29, 14, 23] conduct interactions at the encoding stage by sharing parts of parameter spaces, especially, [23] designs attention modules for feature extraction. While another group of works further studies the parameters sharing scheme in the encoding stage and develops the Branched MTL [27, 36, 15, 2], aiming at manually or automatically determining the task-shared and task-specific branches to minimizing task inconsistencies. In contrast, decoder-focused models, such as PAD-Net [48], employ the multi-model distillation based on initial predictions by deep supervision at the decoding stage, where features of auxiliary tasks are extracted and transferred to target tasks to improve their performance. MTI-Net [38] argues that the information of different tasks only promotes each other at certain scales, and extends the multi-model distillation to the multi-scale fashion. PAP-Net [53] and PSD [55] apply the affinity and the interactions of tasks at pixel and pattern-structure levels respectively. ATRC [3] designs automatically searching for source-to-target task relations by adaptive task-relational context heads with Neural Architecture Search (NAS) techniques. InvPT [50] is the first transformer-based joint learning architecture of global spatial interaction and simultaneous all-task interaction. Although decoder-focused models have recently developed rapidly with a large number of variants, and achieve significant improvements, the interaction of them is only based on task-specific features. Besides, the idea of directly source-to-target task-pair relation modeling is inefficient. To tackle the information loss and high computational cost of these methods, we adopt a brand new bridge-feature-centric interaction strategy, which absorbs the advantages of both encoder- and decoder-focused methods, in order to maintain comprehensive task interaction along with acceptable resource cost.
@@
Attention in MTL: Recently, many MTL architectures [23, 48, 53, 38, 3, 50] have employed attention modules to dynamically select information among different tasks. In [23], attention is used to extract features of different tasks from the shared backbone. [48, 38] distill auxiliary tasks through the attention to extract useful features for target tasks. Since self-attention is capable of capturing global receptive fields by considering the global correlation in feature maps [43, 13], some MTL works directly introduce self-attention into networks [53]. However, these works do not further explore the distribution and interaction of attention among different tasks. With the development of Transformer [39] in vision [10, 42, 45, 25, 47, 5], Multi-head self-attention (MHSA) becomes popular for processing visual scenarios due to its long-range dependencies and context-capture capability. We further study the patterns of MHSA among different tasks and propose the Task Pattern Propagation (TPP) which produces discriminative task-specific features in the early decoding stage to solve the task-pattern-entanglement issue. The transformer-based BFE produces bridge features by global cross-attention between task-generic and task-specific features. Different from simply applying attention to MTL architectures, we focus on structural analysis of task features in the whole process, and introduce priors (e.g. low-level representations, task-specific patterns) with special design accordingly.

### v2

# BridgeNet: Comprehensive and Effective Feature Interactions via Bridge Feature for Multi-task Dense Predictions

@@ Abstract
Multi-task dense prediction aims at handling multiple pixel-wise prediction tasks within a unified network simultaneously for visual scene understanding. However, cross-task feature interactions of current methods are still suffering from incomplete levels of representations, less discriminative semantics in feature participants, and inefficient pair-wise task interaction processes. To tackle these under-explored issues, we propose a novel BridgeNet framework, which extracts comprehensive and discriminative intermediate Bridge Features, and conducts interactions based on them. Specifically, a Task Pattern Propagation (TPP) module is firstly applied to ensure highly semantic task-specific feature participants are prepared for subsequent interactions, and a Bridge Feature Extractor (BFE) is specially designed to selectively integrate both high-level and low-level representations to generate the comprehensive bridge features. Then, instead of conducting heavy pair-wise cross-task interactions, a Task-Feature Refiner (TFR) is developed to efficiently take guidance from bridge features and form final task predictions. To the best of our knowledge, this is the first work considering the completeness and quality of feature participants in cross-task interactions. Extensive experiments are conducted on NYUD-v2, Cityscapes and PASCAL Context benchmarks, and the superior performance shows the proposed architecture is effective and powerful in promoting different dense prediction tasks simultaneously.
@@
Multi-task Learning, Dense Prediction, Representation Learning.
@@ 1 Introduction
Figure 1: Illustration of multi-task interaction strategies. (a) Encoder-focused [30, 24, 15, 28, 37, 16, 2]: the task-specific features directly interact with the task-generic features from the common backbone. The task-generic features contain rich low-level representations but lack high-level representations. The interaction complexity is $O\left(n\right)$. (b) Decoder-focused [49, 10, 3, 55, 57, 51]: the task-specific features are firstly produced by deep supervision, and then interactions are conducted only based on task-specific features, which has discriminative high-level representations, but low-level representations are absent. The pair-wise interaction complexity is $O\left(n^{2}\right)$. (c) BridgeNet (ours): the bridge features are produced from both task-generic and task-specific features and have both rich high-level and low-level representations. The subsequent interactions based on bridge features only have $O\left(n\right)$ complexity. On the right part of the legend, we show some examples of representations from different levels in the feature map: the low-level representations contain less discriminative task information but are rich in details like boundaries, corners and textures, on the contrary, the high-level representations are smooth with less image details, but highly correlated to the task properties, like the highlighted floor with semantic, or areas with larger distance from the depth sensor.
@@
According to where task interactions happen, existing multi-task works are roughly divided into encoder-focused and decoder-focused methods [38]. In particular, the encoder-focused methods conduct interactions of different tasks in the encoding stage, as shown in Fig. 1 (a), task-specific features interact directly with task-generic features produced by the shared backbone, to gain knowledge from multi-task shared representations and thus achieve indirect cross-task interactions. Since the image backbone/encoder is usually initialized with pre-trained parameters that contain abundant visual prior knowledge, and are supervised by gradients from all tasks during the training, it can produce stronger representation with rich low-level representations, i.e. fundamental image details such as edges, corners, textures, etc., as shown in the right legend of Fig. 1. For pixel-level dense prediction tasks like semantic segmentation, these representations are essential in gaining basic pixel relations and forming accurate semantic masks. However, for joint dense predictions among multiple tasks, these low-level representations are task-generic, which means they only contain general image structures, but lack distinguishable semantics that relate to each specific task. Directly conducting interactions based on task-generic features is sub-optimal, since the mining of discriminative task-specific features is insufficient and consequently limits the role of cross-task interaction [38].
@@
To alleviate the above issue, the decoder-focused methods conducting interactions in the decoding stage have been developed, where discriminative high-level representations, i.e. abstract and semantic representations derived from images such as key-points, objects, geometries, etc., is perceived from ground-truth labels to facilitate feature interactions, as shown in the right legend of Fig. 1. Specifically, the initial task predictions are firstly produced by preliminary decoders with deep supervision in the early decoding stage [49, 55, 57, 39, 51], aiming to decouple the generic encoder features and discover task-specific high-level representations according to the corresponding labels. Subsequently, the cross-task interactions are conducted based on high-level representations by modules such as multi-modal distillation [49] or the ATRC [3]. These methods try to model task-pair relations, as shown in Fig. 1 (b), by extracting and transferring useful task knowledge from one task to another based on task-specific features, most of them achieve better performance compared with encoder-focused methods.
@@
i) Firstly, the cross-task interactions are suffering from incomplete levels of representations in the participants, i.e. either task-generic features with low-level representations or task-specific features with high-level representations are exploited in prior works, which leads to incomplete participants for feature interaction. Though decoder-focused methods have validated that the interactions with discriminative task-specific features are more effective, in which high-level representations are gained in the initial decoding stage. However, the essential low-level representations with abundant basic image details gradually disappear since they are not directly required by supervision signals, and subsequently, in the following task interaction stage, only high-level representations are involved while low-level representations that produce essential dense prediction priors among different tasks are absent, leading to incomplete task interactions and limiting the model performance.
ii) Secondly, the cross-task interactions are also suffering from less discriminative semantics in the participants. We observe that the task-specific features produced by the preliminary decoders are usually of low quality with less discriminative representations (as shown in Fig 3 right), which negatively affects the subsequent interactions based on them. By analyzing the task patterns, which are defined as the attention distributions or feature structures specifically related to a certain task (as shown in Fig. 3 left), we discovered that task patterns significantly differ on different tasks. While in the task-generic features produced by the shared encoder, patterns are usually highly entangled, which makes it difficult to decompose each specific task pattern, and simply imposing supervision on preliminary decoders is not able to fundamentally solve this task-pattern entanglement issue. Thus a more effective way of disentangling meaningful and semantically high-level representations from the shared task-generic features is required for producing high-quality interaction participants.
@@
iii) The cross-task interaction way is inefficient with high extension costs. Here exemplified by Fig. 1 (b). Since previous decoder-focused models need to consider pair-wise task relations [49, 39, 3, 51], with the task number increasing, the complexity of task interactions will increase in $O\left(n^{2}\right)$, which is costly and limits the extension of previous methods on the settings with more tasks.
@@
Figure 2: The overview of our proposed method. We take depth estimation and semantic segmentation as examples. We use a shared image backbone to encode task-generic features, and a set of preliminary decoders with deep supervision are used for task-specific feature extraction. Different from decoder-focused methods, a TPP is added before the initial predictions are formed to tackle the task-pattern-entanglement issue. The produced high-quality task-specific features are not used for interaction directly, they are firstly processed by BFE at each scale to gain low-level representations and form the bridge features. Then, the multi-scale bridge features are used to gradually refine task-specific features by TFR. The outputs of TFR are aggregated and fed into task-specific heads for final predictions. The detailed structures of TPP (Sec. 3.2, Fig. 4 (a)), BFE (Sec. 3.3, Fig. 5) and TFR (Sec. 3.4, Fig. 4 (b)) will be given later.
@@
To this end, we focus on improving the interaction quality for superior Multi-task Learning from the above three aspects, and propose a novel BridgeNet framework, which is based on the interaction with a comprehensive intermediate feature, namely the Bridge Feature. As shown in Fig. 1 (c), to solve the first challenge, we propose to extract and take advantage of the bridge feature which contains both rich low-level and high-level representations to ensure the integrity of feature interactions. To solve the second challenge, we propose to disentangle task patterns by jointly learning and distributing task patterns to corresponding tasks, in order to produce high-quality task-specific features serving as interaction participants. And to solve the third challenge, we propose to conduct interactions directly between the bridge features and each task-specific feature, which only involves $O\left(n\right)$ complexity of task interactions. Specifically, our proposed method consists of a shared encoder backbone for the task-generic features production, the preliminary decoders with Task Pattern Propagation (TPP) in the early decoding stage producing high-quality task-specific features and handling the task-pattern-entanglement issue. The specially designed Bridge Feature Extractor (BFE), which has a transformer-based structure with cross-attention. By globally querying task-specific features with the task-generic features at each scale, the BFE selects high-level representations correlated to every task and injects them into the task-generic features with rich low-level representations. Thus the extracted powerful bridge features contain comprehensive task representations in both high-level and low-level, and fertilize the formation of final predictions by the Task Feature Refiner (TFR). The overall bridge-feature-based multi-task framework is named as BridgeNet.
@@
    We discover that incomplete, low-quality interaction participants and inefficient interaction processes exist in current cross-task interaction designs of multi-task dense prediction frameworks. To tackle these issues, we propose the novel BridgeNet framework, which performs a comprehensive interaction based on the bridge features containing both low-level and high-level representations. To the best of our knowledge, bridge features with comprehensive types of representations are involved in the multi-task dense prediction for the first time, which are exploited to transfer cross-task knowledge and fertilize the final task predictions.
@@
    A Task Pattern Propagation (TPP) module is firstly applied to avoid the task-pattern-entanglement issue and prepare high-quality participants for the subsequent interactions. A transformer-based Bridge Feature Extractor (BFE) is designed to extract bridge features from both task-generic and task-specific features for comprehensive interactions. Finally, a Task Feature Refiner (TFR) is applied to take advantage of bridge features and refine the final task predictions.
@@
Figure 3: Left: Visualization of different patterns of two tasks (semantic segmentation and depth estimation). The semantic feature focuses on objects with various semantic information, like cabinets and floors, while depth attention focuses more on the surface and edge features with geometry information. Right: Visualization of the task-specific features with (w.) and without (w.o.) TPP. The task-generic features produced by the shared encoder show the task-pattern-entanglement issue, which is significantly different from the distributions of task labels, and the contained patterns are implicit and ambiguous. Without TPP, the decomposed task-specific features serving as the interaction participants are still struggling with the lack of discriminative semantics, which negatively affects the subsequent interaction process (BFE, TFR) and eventually produces low-quality interaction outputs. However, with TPP, the task-specific features can obtain well-decomposed representations that have more similar distributions to their ground-truth labels (like the highlighted floor area in semantic segmentation and door area in depth estimation), and thus boosts the subsequent interaction process to produce high-quality discriminative features.
@@ 2 Related Works
Multi-Task Learning Architectures: With the rapid development of deep CNNs, a lot of MTL works have achieved promising results [30, 15, 24, 49, 54, 55, 57, 28, 37, 39, 2, 34, 16, 21, 52, 47, 18, 51]. As mentioned in Sec. 1, existing MTL works are roughly divided into encoder-focused and decoder-focused architectures. In addition, the sharing ways of network parameters are categorized into hard and soft parameter sharing [38, 10], which correspond to directly shared parameters or indirectly constrained by regularization or losses, respectively. The encoder-focused models in [15, 34, 30] have specifically designed shared structures to explore backbone task-generic features in the encoding stage, while the decoder-focused models in [38] adopt hard parameters sharing in the encoding stage to reduce redundant parameters and FLOPs, and specially focus on the interaction among different task-specific features in decoding stages. In this paper, our proposed method is based on the decoder-focused architecture, however, also takes advantage of the encoder-focused models, i.e. the exploring of task-generic features with rich low-level representations, and computational-friendly multi-task interaction ways.
@@
Task Interaction Strategies in MTL: In MTL, the learned representation from one task may be beneficial for other tasks, so it is essential to design task interaction manners to promote mutual performances, and avoid potential information inconsistencies. In the field of knowledge distillation, a multi-task student network is distilled through several single-task teacher networks, which transfer multi-task knowledge to the student network and achieve good performance [52, 21]. However, this network requires to pretrain a group of teacher networks which costs a lot by increasing the number of tasks. The typical encoder-focused models [30, 15, 24] conduct interactions at the encoding stage by sharing parts of parameter spaces, especially, [24] designs attention modules for feature extraction. While another group of works further studies the parameters sharing scheme in the encoding stage and develops the Branched MTL [28, 37, 16, 2], aiming at manually or automatically determining the task-shared and task-specific branches to minimizing task inconsistencies. In contrast, decoder-focused models, such as PAD-Net [49], employ the multi-model distillation based on initial predictions by deep supervision at the decoding stage, where features of auxiliary tasks are extracted and transferred to target tasks to improve their performance. MTI-Net [39] argues that the information of different tasks only promotes each other at certain scales, and extends the multi-model distillation to the multi-scale fashion. PAP-Net [55] and PSD [57] apply the affinity and the interactions of tasks at pixel and pattern-structure levels respectively. ATRC [3] designs automatically searching for source-to-target task relations by adaptive task-relational context heads with Neural Architecture Search (NAS) techniques. InvPT [51] is the first transformer-based joint learning architecture of global spatial interaction and simultaneous all-task interaction. Although decoder-focused models have recently developed rapidly with a large number of variants, and achieve significant improvements, the interaction of them is only based on task-specific features. Besides, the idea of directly source-to-target task-pair relation modeling is inefficient. To tackle the information loss and high computational cost of these methods, we adopt a brand new BridgeNet architecture, which absorbs the advantages of both encoder- and decoder-focused methods, in order to maintain comprehensive task interaction along with acceptable resource cost.
@@
Attention in MTL: Recently, many MTL architectures [24, 49, 55, 39, 3, 51, 53] have employed attention modules to dynamically select information among different tasks. In [24], attention is used to extract features of different tasks from the shared backbone. [49, 39] distill auxiliary tasks through the attention to extract useful features for target tasks. Since self-attention is capable of capturing global receptive fields by considering the global correlation in feature maps [44, 14], some MTL works directly introduce self-attention into networks [55]. However, these works do not further explore the distribution and interaction of attention among different tasks. With the development of Transformer [40] in vision [11, 43, 46, 26, 48, 5], Multi-head self-attention (MHSA) becomes popular for processing visual scenarios due to its long-range dependencies and context-capture capability. We further study the patterns of MHSA among different tasks and propose the Task Pattern Propagation (TPP) which produces discriminative task-specific features in the early decoding stage to solve the task-pattern-entanglement issue. The transformer-based BFE produces bridge features by global cross-attention between task-generic and task-specific features. Different from simply applying attention to MTL architectures, we focus on structural analysis of task features in the whole interaction process, and introduce priors (e.g. low-level image representations, task patterns) with special design accordingly.